OpenStudy (anonymous):

in the question: -2(3x+4) do you multiply everything by -2 or just 2?

mathslover (mathslover):

a(b+c) = ab + ac Similarly : -2(3x+4) = -2(3x) -2(4)

OpenStudy (anonymous):

by -2